At 2000°C the equilibrium constant for the reaction is Kc = 2.4 103. 2 NO(g) N2(g) + O2(g) If the initial concentration of NO is 0.180 M, what are the equilibrium concentrations of NO, NO is not 0.002, 0.004, 0.02. 0.00180 NO is IMPORTANT, need the CORRECT answer
Answer #1

ICE Table:

                    [NO]                [N2]                [O2]              


initial             0.18                0                   0                 

change              -2x                 +1x                 +1x               

equilibrium         0.18-2x             +1x                 +1x               

Equilibrium constant expression is
Kc = [N2]*[O2]/[NO]^2
2400.0 = (1*x)^2/(0.18-2*x)^2
sqrt(2400.0) = (1*x)/(0.18-2*x)
48.99 = (1*x)/(0.18-2*x)
8.81816-97.97959*x = 1*x
8.81816-98.97959*x = 0
x = 0.0891

At equilibrium:
[NO] = 0.180-2x = 0.180-2*0.0891 = 0.00182 M

Answer: 0.00182 M